OpenStudy (anonymous):

Please help with this calculus problem! For what values of x is the graph of e^-x62 concave downwards. Explain all steps in detail please.

OpenStudy (callisto):

\[e^{-x62} \] or \[e^{-x^{62}} \] or ???

OpenStudy (anonymous):

differentiate your function twice (i'm assuming you can differentiate since you're in calculus) f'' = 3844e^-64x, which is positive for all x, this means that since f''>0 for all real x, the original function f is concave up for all x, therefore it never decreases

OpenStudy (anonymous):

The way I wrote it, is the exact way the question wrote it

OpenStudy (anonymous):

sorry, by decreases i mean concave downwards

OpenStudy (anonymous):

Ok thanks, is that the answer to the question?

OpenStudy (anonymous):

assuming your function is f = {e to the power of -62 times x} then yes

OpenStudy (anonymous):

So how would it be written as, \[e ^{-x62}\]?

OpenStudy (anonymous):

And can you show the steps you took to get the second derivative, I dont know how to find derivatives

OpenStudy (callisto):

Assume your question is \[e^{-x62}\] First derivative\[d/dx (e^{-x62}) = d/dx (e^{-62x})= d/dx (e^{-62x}) =-62e^{-62x}\] second derivatives \[d/dx (-62e^{-62x}) = (-62)(-62)e^{-62x}\]

OpenStudy (anonymous):

So after we have the second derivative what do we do?

OpenStudy (callisto):

I think... put second derivative =0 ... Not sure for this case... :(

OpenStudy (anonymous):

refer to my answer above ^

OpenStudy (anonymous):

Is the second derivative 3844e^-62x?

OpenStudy (anonymous):

K thanks anyway callisto :) I really appreciate it

OpenStudy (callisto):

Is the second derivative 3844e^-62x? -> yes